factor the following trinomial x^2+8xb+12b^2
1
Expert's answer
2011-10-18T12:29:19-0400
let's find roots of the following equation
x^2+8xb+12b^2=0
Descriminant D=(64-48)b^2=(4b)^2
Hence x1=(-8b+4b)/2=-2b
x2=(-8b-4b)/2=-6b
So we have x^2+8xb+12b^2=(x+2b)(x+6b)